# Adversarial Domain Adaptation for Cross-user Activity Recognition Using Diffusion-based Noise-centred Learning

### v1

@@ V Conclusion
In this paper, we have presented a novel method Diff-Noise-Adv-DA for cross-user HAR domain adaptation, leveraging diffusion models with adversarial regularization. Our approach enhances the forward diffusion process by incorporating classification information and improves the reverse denoising process by introducing adversarial regularization. The training objectives are designed to maximize data likelihood while minimizing domain discrepancy, promoting better alignment between source and target distributions. Experimental results demonstrate the effectiveness of our method in achieving robust cross-user HAR domain adaptation, highlighting its potential for real-world applications.

### v2

@@ V Conclusion
In this paper, we introduced DNA-DA, a novel method for cross-user HAR domain adaptation. Our approach utilizes diffusion models combined with adversarial learning for domain adaptation. Specifically, we enhance the forward diffusion process by incorporating classification information into the noise, and we improve the reverse denoising process through adversarial regularization. The training objectives are set to achieve a max-min adversarial learning goal, which helps in better aligning source and target users’ data distributions. Experimental results demonstrate the effectiveness of our method in achieving robust cross-user HAR domain adaptation, showcasing its potential for real-world applications.
For future work, we plan to further test the DNA-DA method with activities that have more distinct data distributions between users. This will help us evaluate the method’s versatility and robustness in handling more challenging tasks. Additionally, we intend to broaden our experimental approach by including a wider range of user groups, such as individuals with different physical characteristics and those from different hand usage habits(i.e. left-handed or right-handed activities). This will help us determine how well DNA-DA performs across various segments of the population.
To strengthen our evaluation, we are developing our own datasets that cover a wide range of real-world scenarios. These custom datasets will provide insights into how well DNA-DA adapts to different situations beyond the controlled lab environment. This effort aims not only to validate our approach but also to contribute new and varied datasets to the research community.